OpenStudy (tabithax3):

Solve the equation in the interval [0, 2π) cos2x=√2-cos2x

OpenStudy (raden):

move -cos2x to left side, what u get

OpenStudy (raden):

ur question, is cos2x=√2 - cos2x or cos2x=√(2-cos2x)

OpenStudy (tabithax3):

@RadEn it's cos2x=√2 - cos2x. and for the first question it's 2cos2x

OpenStudy (raden):

ok, it can be 2cos2x = √2 or cos2x = 1/2 * √2 , right ?

OpenStudy (tabithax3):

@RadEn yes. that's what i did, but i stopped there

OpenStudy (raden):

just remember, what's angles when cos of that angle has the value 1/2 * √2 ?

OpenStudy (raden):

it's the special degrees

OpenStudy (tabithax3):

π/4

OpenStudy (raden):

any else?

OpenStudy (raden):

pi/4 just in the first quadrant, but in the 4th quadrant is .... degreess

OpenStudy (tabithax3):

and 7π/4

OpenStudy (raden):

yes, correct

OpenStudy (raden):

so, the equation can be : case I : cos2x = 1/2 * √2 cos2x = cospi/4 cos2x = cos(2pi*n +- pi/4) so, 2x = 2pi*n +- pi/4 or x = (2pi*n +- pi/4)/2 = pi*n +- pi/8 (this is just general solution) do same idea, for case II : cos2x = 1/2 * √2 cos2x = cos7pi/4 continue it....

OpenStudy (tabithax3):

the answers will be {π/8, 9π/8, 7π/8, 15π/8}

OpenStudy (raden):

yes, for the 2nd case happened duplicates for solution from the 1st case u are right

OpenStudy (tabithax3):

i understand it now. thank you so much @RadEn :)

OpenStudy (raden):

very welcome :)